# Exact overlaps for ”all” integrable matrix product states of rational spin chains

### v1

@@ ACKNOWLEDGMENTS
$|1\rangle\equiv Z,\quad|2\rangle\equiv Y,\quad|3\rangle\equiv X,\quad|4\rangle\equiv\bar{X},\quad|5\rangle\equiv\bar{Y},\quad|6\rangle\equiv\bar{Z}.$ (S-71)
Our MPS is built from a one-side operator
$\bar{\omega}=\left\{S_{1},S_{2},S_{3},S_{3},S_{2},S_{1}\right\}$ (S-72)
as
$\langle MPS|=\sum_{i_{1},\dots,i_{J}}\mathrm{Tr}\left[\omega_{i_{J}}\dots\omega_{i_{1}}\right]\langle i_{1},i_{2},\dots,i_{J-1},i_{J}|.$ (S-73)
We defined the spin operators $S_{1},S_{2},S_{3}$ which satisfies the $\mathfrak{su}_{2}$ algebra
$\left[S_{j},S_{k}\right]=\sum_{l=1}^{3}i\epsilon_{j,k,l}S_{l},$ (S-74)
We will also need the $\mathfrak{su}_{2}$ Casimir
$S^{2}=S_{1}^{2}+S_{2}^{2}+S_{3}^{2}.$ (S-75)
The first step to apply our formula is to get monodromy matrices. Since the $\mathfrak{sl}_{4}\cong\mathfrak{o}_{6}$ we could use the $\mathfrak{gl}_{4}$ or $\mathfrak{o}_{6}$ formulas. Now we continue with the $\mathfrak{gl}_{4}$ convention. The $\mathfrak{gl}_{4}$ monodromy matrices can be defined in the following way. Let $e_{i}$ be the usual basis in $\mathbb{C}^{4}$. Let $e_{i,j}$ be the unit matrices in $\mathbb{C}^{4}$ as $e_{i,j}e_{k}=\delta_{j,k}e_{i}$. Let $e_{[i,j]}$ for $1\leq i<j\leq 4$ be a basis of $\mathbb{C}^{6}$ which coming from the anti-symmetrization of $\mathbb{C}^{4}\otimes\mathbb{C}^{4}$. Let $e_{[i,j],[k,l]}$ be the unit matrices in $\mathbb{C}^{6}$ as $e_{[i,j],[k,l]}e_{[a,b]}=\delta_{k,a}\delta_{l,b}e_{[i,j]}$. We can identify the six dimensional basis as
@@
$\mathcal{L}(u)=\sum_{i,j=1}^{4}e_{i,j}\otimes\left(\delta_{i,j}+\frac{1}{u+1/2}\mathcal{E}_{j,i}\right),\qquad\widehat{\mathcal{L}}(u)=\sum_{i,j=1}^{4}e_{i,j}\otimes\left(\delta_{i,j}+\frac{1}{-u+1/2}\mathcal{E}_{i,j}\right)$ (S-78)
@@
$\widehat{T}_{0}(u)$ =^$=\widehat{\mathcal{L}}_{0,J}(u)\dots\widehat{\mathcal{L}}_{0,2}(u)\widehat{\mathcal{L}}_{0,1}(u).$ (S-79)
@@
The next step is to find the $K$-matrix. The $K$-matrix is the solution of the $KT$-relation for $J=1$
@@
Next step is to calculate the $G$-operators which are defined in (S-III.1). We need the nested $K$-matrices $\mathbf{K}^{(2)},\mathbf{K}^{(3)}$. Since $\mathbf{K}_{1,1}=1$ the second $K$-matrix is simple defined as
@@
$\mathbf{K}^{(2)}=\left(\begin{array}[]{ccc}\mathbf{K}_{2,2}^{(2)}&\mathbf{K}_{2,3}^{(2)}&\mathbf{K}_{2,4}^{(2)}\\ \mathbf{K}_{3,2}^{(2)}&\mathbf{K}_{3,3}^{(2)}&\mathbf{K}_{3,4}^{(2)}\\ \mathbf{K}_{4,2}^{(2)}&\mathbf{K}_{4,3}^{(2)}&\mathbf{K}_{4,4}^{(2)}\end{array}\right)=\left(\begin{array}[]{ccc}1-\frac{4S_{1}^{2}}{(2u+1)^{2}}&\frac{2i}{2u+1}S_{3}-\frac{4S_{1}.S_{2}}{(2u+1)^{2}}&\frac{-2i}{2u+1}S_{2}-\frac{4S_{1}.S_{3}}{(2u+1)^{2}}\\ \frac{-2i}{2u+1}S_{3}-\frac{4S_{2}.S_{1}}{(2u+1)^{2}}&1-\frac{4S_{2}^{2}}{(2u+1)^{2}}&\frac{2i}{2u+1}S_{1}-\frac{4S_{2}.S_{3}}{(2u+1)^{2}}\\ \frac{2i}{2u+1}S_{2}-\frac{4S_{3}.S_{1}}{(2u+1)^{2}}&\frac{-2i}{2u+1}S_{1}-\frac{4S_{3}.S_{2}}{(2u+1)^{2}}&1-\frac{4S_{3}^{2}}{(2u+1)^{2}}\end{array}\right).$ (S-83)
@@
$\mathbf{K}_{2,2}^{(2)}(u+1)\mathbf{K}_{a,2}^{(2)}(u)=\left(\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)-\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\right)\mathbf{K}_{2,2}^{(2)}(u)$ (S-85)
@@
$\mathbf{K}_{a,2}^{(2)}(u)\left[\mathbf{K}_{2,2}^{(2)}(u)\right]^{-1}=\left[\mathbf{K}_{2,2}^{(2)}(u+1)\right]^{-1}\left(\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)-\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\right)$ (S-86)
@@
$\mathbf{K}_{a,b}^{(3)}(u)=\left[\mathbf{K}_{2,2}^{(2)}(u+1)\right]^{-1}\tilde{\mathbf{K}}_{a,b}^{(3)}(u)$
where we introduced $2\times 2$ matrix without inversion
$\tilde{\mathbf{K}}_{a,b}^{(3)}(u)=\mathbf{K}_{2,2}^{(2)}(u+1)\mathbf{K}_{a,b}^{(2)}(u)-\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)\mathbf{K}_{2,b}^{(2)}(u)+\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\mathbf{K}_{2,b}^{(2)}(u).$ (S-87)
@@
$\tilde{\mathbf{K}}^{(3)}=\left(\begin{array}[]{cc}\mathbf{K}_{3,3}^{(2)}&\mathbf{K}_{3,4}^{(2)}\\ \mathbf{K}_{4,3}^{(2)}&\mathbf{K}_{4,4}^{(2)}\end{array}\right)=\frac{(2u+1)(2u+3)-4S^{2}}{(2u+1)(2u+3)}\left(\begin{array}[]{cc}1&\frac{2i}{2u+3}S_{1}\\ \frac{-2i}{2u+3}S_{1}&1\end{array}\right),$ (S-88)
@@
$\begin{split}\mathbf{G}^{(1)}(u)=&\mathbf{K}_{1,1}(u)=1\\ \mathbf{G}^{(2)}(u)=&\mathbf{K}_{2,2}^{(2)}(u)=\frac{(2u+1)^{2}-4S_{1}^{2}}{(2u+1)^{2}}\\ \mathbf{G}^{(3)}(u)=&\mathbf{K}_{3,3}^{(3)}(u)=\frac{(2u+3)}{(2u+1)}\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)^{2}-4S_{1}^{2}}\\ \mathbf{G}^{(4)}(u)=&\mathbf{K}_{4,4}^{(4)}(u)=\mathbf{K}_{4,4}^{(3)}(u)-\mathbf{K}_{4,3}^{(3)}(u)\left[\mathbf{K}_{3,3}^{(3)}(u)\right]^{-1}\mathbf{K}_{3,4}^{(3)}(u)\\ =&\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)(2u+1)}.\end{split}$ (S-89)
@@
$\begin{split}\mathbf{F}^{(1)}(u)=&\left[\mathbf{G}^{(1)}(u)\right]^{-1}\mathbf{G}^{(2)}(u)=\frac{(2u+1)^{2}-4S_{1}^{2}}{(2u+1)^{2}}\\ \mathbf{F}^{(2)}(u)=&\left[\mathbf{G}^{(2)}(u)\right]^{-1}\mathbf{G}^{(3)}(u)=\frac{(2u+1)(2u+3)}{(2u+1)^{2}-4S_{1}^{2}}\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)^{2}-4S_{1}^{2}}\\ \mathbf{F}^{(3)}(u)=&\left[\mathbf{G}^{(3)}(u)\right]^{-1}\mathbf{G}^{(4)}(u)=\frac{(2u+3)^{2}-4S_{1}^{2}}{(2u+3)^{2}}\end{split}$ (S-90)
@@
$\begin{split}\mathcal{F}_{m}^{(1)}(u)&=\frac{(2u+1-2m)(2u+1+2m)}{(2u+1)^{2}},\\ \mathcal{F}_{m}^{(2)}(u)=&\frac{(2u+1)(2u+3)}{(2u+1-2m)(2u+1+2m)}\frac{(2u+1-2s)(2u+3+2s)}{(2u+3-2m)(2u+3+2m)},\\ \mathcal{F}_{m}^{(3)}(u)=&\frac{(2u+3-2m)(2u+3+2m)}{(2u+3)^{2}}.\end{split}$ (S-92)

### v2

@@ Acknowledgments
$|1\rangle\equiv Z,\quad|2\rangle\equiv Y,\quad|3\rangle\equiv X,\quad|4\rangle\equiv\bar{X},\quad|5\rangle\equiv\bar{Y},\quad|6\rangle\equiv\bar{Z}.$ (S-72)
The zeroth step to apply our formula is to get monodromy matrices. Since the $\mathfrak{sl}_{4}\cong\mathfrak{o}_{6}$ we could use the $\mathfrak{gl}_{4}$ or $\mathfrak{o}_{6}$ formulas. Now we continue with the $\mathfrak{gl}_{4}$ convention. The $\mathfrak{gl}_{4}$ monodromy matrices can be defined in the following way. Let $e_{i}$ be the usual basis in $\mathbb{C}^{4}$. Let $e_{i,j}$ be the unit matrices in $\mathbb{C}^{4}$ as $e_{i,j}e_{k}=\delta_{j,k}e_{i}$. Let $e_{[i,j]}$ for $1\leq i<j\leq 4$ be a basis of $\mathbb{C}^{6}$ which coming from the anti-symmetrization of $\mathbb{C}^{4}\otimes\mathbb{C}^{4}$. Let $e_{[i,j],[k,l]}$ be the unit matrices in $\mathbb{C}^{6}$ as $e_{[i,j],[k,l]}e_{[a,b]}=\delta_{k,a}\delta_{l,b}e_{[i,j]}$. We can identify the six dimensional basis as
@@
$\mathcal{L}(u)=\sum_{i,j=1}^{4}e_{i,j}\otimes\left(\delta_{i,j}+\frac{1}{u+1/2}\mathcal{E}_{j,i}\right),\qquad\widehat{\mathcal{L}}(u)=\sum_{i,j=1}^{4}e_{i,j}\otimes\left(\delta_{i,j}+\frac{1}{-u+1/2}\mathcal{E}_{i,j}\right).$ (S-75)
@@
$\widehat{T}_{0}(u)$ =^$=\widehat{\mathcal{L}}_{0,J}(u)\dots\widehat{\mathcal{L}}_{0,2}(u)\widehat{\mathcal{L}}_{0,1}(u).$ (S-76)
@@
S-VII.1.1 $\mathfrak{o}_{4}=\mathfrak{o}_{3}\oplus\mathfrak{o}_{3}$ symmetric boundary states
Our MPS is built from a one-side operator
$\begin{split}\bar{\omega}&=\left\{S_{1},S_{2},S_{3},S_{3},S_{2},S_{1}\right\},\\ \langle\mathrm{MPS}|&=\sum_{i_{1},\dots,i_{J}}\mathrm{Tr}\left[\omega_{i_{J}}\dots\omega_{i_{1}}\right]\langle i_{1},i_{2},\dots,i_{J-1},i_{J}|.\end{split}$ (S-77)
We defined the spin operators $S_{1},S_{2},S_{3}$ which satisfy the $\mathfrak{su}_{2}$ algebra
$\left[S_{j},S_{k}\right]=\sum_{l=1}^{3}i\epsilon_{j,k,l}S_{l}.$ (S-78)
We will also need the $\mathfrak{su}_{2}$ Casimir operator
$S^{2}=S_{1}^{2}+S_{2}^{2}+S_{3}^{2}.$ (S-79)
The first step is to find the $K$-matrix. The $K$-matrix is the solution of the $KT$-relation for $J=1$
@@
Next step is to calculate the $G$-operators which are defined in S-III.1. We need the nested $K$-matrices $\mathbf{K}^{(2)},\mathbf{K}^{(3)}$. Since $\mathbf{K}_{1,1}=1$ the second $K$-matrix is simply defined as
@@
$\mathbf{K}^{(2)}=\left(\begin{array}[]{ccc}\mathbf{K}_{2,2}^{(2)}&\mathbf{K}_{2,3}^{(2)}&\mathbf{K}_{2,4}^{(2)}\\ \mathbf{K}_{3,2}^{(2)}&\mathbf{K}_{3,3}^{(2)}&\mathbf{K}_{3,4}^{(2)}\\ \mathbf{K}_{4,2}^{(2)}&\mathbf{K}_{4,3}^{(2)}&\mathbf{K}_{4,4}^{(2)}\end{array}\right)=\left(\begin{array}[]{ccc}1-\frac{4S_{1}^{2}}{(2u+1)^{2}}&\frac{2i}{2u+1}S_{3}-\frac{4S_{1}S_{2}}{(2u+1)^{2}}&\frac{-2i}{2u+1}S_{2}-\frac{4S_{1}S_{3}}{(2u+1)^{2}}\\ \frac{-2i}{2u+1}S_{3}-\frac{4S_{2}S_{1}}{(2u+1)^{2}}&1-\frac{4S_{2}^{2}}{(2u+1)^{2}}&\frac{2i}{2u+1}S_{1}-\frac{4S_{2}S_{3}}{(2u+1)^{2}}\\ \frac{2i}{2u+1}S_{2}-\frac{4S_{3}S_{1}}{(2u+1)^{2}}&\frac{-2i}{2u+1}S_{1}-\frac{4S_{3}S_{2}}{(2u+1)^{2}}&1-\frac{4S_{3}^{2}}{(2u+1)^{2}}\end{array}\right).$ (S-83)
@@
$\mathbf{K}_{2,2}^{(2)}(u+1)\mathbf{K}_{a,2}^{(2)}(u)=\left(\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)-\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\right)\mathbf{K}_{2,2}^{(2)}(u),$ (S-85)
@@
$\mathbf{K}_{a,2}^{(2)}(u)\left[\mathbf{K}_{2,2}^{(2)}(u)\right]^{-1}=\left[\mathbf{K}_{2,2}^{(2)}(u+1)\right]^{-1}\left(\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)-\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\right).$ (S-86)
@@
$\mathbf{K}_{a,b}^{(3)}(u)=\left[\mathbf{K}_{2,2}^{(2)}(u+1)\right]^{-1}\tilde{\mathbf{K}}_{a,b}^{(3)}(u),$ (S-87)
where we introduced a $2\times 2$ matrix without inversion:
$\tilde{\mathbf{K}}_{a,b}^{(3)}(u)=\mathbf{K}_{2,2}^{(2)}(u+1)\mathbf{K}_{a,b}^{(2)}(u)-\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)\mathbf{K}_{2,b}^{(2)}(u)+\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\mathbf{K}_{2,b}^{(2)}(u).$ (S-88)
@@
$\tilde{\mathbf{K}}^{(3)}=\left(\begin{array}[]{cc}\tilde{\mathbf{K}}_{3,3}^{(3)}&\tilde{\mathbf{K}}_{3,4}^{(3)}\\ \tilde{\mathbf{K}}_{4,3}^{(3)}&\tilde{\mathbf{K}}_{4,4}^{(3)}\end{array}\right)=\frac{(2u+1)(2u+3)-4S^{2}}{(2u+1)(2u+3)}\left(\begin{array}[]{cc}1&\frac{2i}{2u+3}S_{1}\\ \frac{-2i}{2u+3}S_{1}&1\end{array}\right).$ (S-89)
@@
$\begin{split}\mathbf{G}^{(1)}(u)=&\mathbf{K}_{1,1}(u)=1,\\ \mathbf{G}^{(2)}(u)=&\mathbf{K}_{2,2}^{(2)}(u)=\frac{(2u+1)^{2}-4S_{1}^{2}}{(2u+1)^{2}},\\ \mathbf{G}^{(3)}(u)=&\mathbf{K}_{3,3}^{(3)}(u)=\frac{(2u+3)}{(2u+1)}\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)^{2}-4S_{1}^{2}},\\ \mathbf{G}^{(4)}(u)=&\mathbf{K}_{4,4}^{(4)}(u)=\mathbf{K}_{4,4}^{(3)}(u)-\mathbf{K}_{4,3}^{(3)}(u)\left[\mathbf{K}_{3,3}^{(3)}(u)\right]^{-1}\mathbf{K}_{3,4}^{(3)}(u)\\ =&\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)(2u+1)}.\end{split}$ (S-90)
@@
$\begin{split}\mathbf{F}^{(1)}(u)=&\left[\mathbf{G}^{(1)}(u)\right]^{-1}\mathbf{G}^{(2)}(u)=\frac{(2u+1)^{2}-4S_{1}^{2}}{(2u+1)^{2}},\\ \mathbf{F}^{(2)}(u)=&\left[\mathbf{G}^{(2)}(u)\right]^{-1}\mathbf{G}^{(3)}(u)=\frac{(2u+1)(2u+3)}{(2u+1)^{2}-4S_{1}^{2}}\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)^{2}-4S_{1}^{2}}\\ \mathbf{F}^{(3)}(u)=&\left[\mathbf{G}^{(3)}(u)\right]^{-1}\mathbf{G}^{(4)}(u)=\frac{(2u+3)^{2}-4S_{1}^{2}}{(2u+3)^{2}},\end{split},$ (S-91)
@@
$\begin{split}\mathcal{F}_{m}^{(1)}(u)&=\frac{(2u+1-2m)(2u+1+2m)}{(2u+1)^{2}},\\ \mathcal{F}_{m}^{(2)}(u)=&\frac{(2u+1)(2u+3)}{(2u+1-2m)(2u+1+2m)}\frac{(2u+1-2s)(2u+3+2s)}{(2u+3-2m)(2u+3+2m)},\\ \mathcal{F}_{m}^{(3)}(u)=&\frac{(2u+3-2m)(2u+3+2m)}{(2u+3)^{2}}.\end{split}$ (S-93)